# Automorphism groups of affine varieties and their Lie algebras

### v1

@@ Abstract.
This is a brief overview of a few selected chapters on automorphism groups of affine varieties. It includes some open questions.
@@ 1. Introduction
We provide a brief overview of a few selected chapters on automorphism groups of affine varieties. It completes existing surveys on the subject, see e.g. [Arz23], [Fre17], [FuKr18], [Kal09], [KPZ17], [Kra89], [Kra96], [Kra17], [Lam24], [Miy24], [Sno89], [vdE00], [vdEKC21], [Wri17], etc.
Throughout the text, $X$ stands for an irreducible affine variety defined over an algebraically closed field ${\mathbb{K}}$ of characteristic zero, ${\mathbb{A}}^{n}$ stands for the affine $n$-space over ${\mathbb{K}}$ and $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$ resp. $\operatorname{{\mathbb{G}}_{\mathrm{m}}}$ stands for the additive resp. multiplicative group of the field ${\mathbb{K}}$.
Already in the case of a smooth affine surface $X$, the group $\operatorname{Aut}(X)$ is often infinite-dimensional and has a rich algebraic structure, see the survey article [KPZ17]. The present survey is focused on various aspects of $\operatorname{Aut}(X)$ as an ind-group, on its algebraically generated subgroups and their Lie algebras. It contains a list of open problems partially borrowed in [FuKr18] and [KrZa22].
The group of regular automorphisms $\operatorname{Aut}(X)$ is an affine ind-group, see e.g. [FuKr18, Theorem 5.1.1]. Recall that an ind-group structure on $\operatorname{Aut}(X)$ is a filtration $\operatorname{Aut}(X)=\bigcup_{i=1}^{\infty}A_{i}$ where
$\ldots\subseteq A_{i}\subseteq A_{i+1}\subseteq\ldots$
is an increasing sequence of affine varieties and their closed embeddings such that for each $i,j$ and some $n(i,j)$ and $m(i)$ the action $A_{i}\times X\to X$, the multiplication $A_{i}\times A_{j}\to A_{n(i,j)}$ and the inversion $A_{i}\to A_{m(i)}$ are morphisms of affine varieties, see [FuKr18].
The ind-group $\operatorname{Aut}(X)$ carries a Zariski ind-topology. A subset $S\subseteq\operatorname{Aut}(X)$ is closed if $S\cap A_{i}$ is a Zariski closed subset of $A_{i}$ for each $i$. A closed algebraic subset of $A_{i}$ defines an algebraic family of automorphisms of $X$; such families were studied in [Ram64].
A faithful action of a linear algebraic group $G$ on $X$ defines an embedding of $G$ onto an algebraic subgroup of $\operatorname{Aut}(X)$. Any such subgroup is a closed algebraic subset of some $A_{i}$.
@@
Let $\mathfrak{G}=\operatorname{Aut}(X)$ and let $\operatorname{Vec}(X)$ stand for the Lie algebra of regular vector fields on $X$. Recall that $\operatorname{Vec}(X)$ is naturally isomorphic to the Lie algebra $\operatorname{Der}(\mathcal{O}(X))$ of derivations of the ${\mathbb{K}}$-algebra $\mathcal{O}(X)$ of regular functions on $X$. In turn, $\operatorname{Lie}(\operatorname{Aut}(X))$ can be naturally identified with a subalgebra $\operatorname{Vec}^{\mathrm{c}}({\mathbb{A}}^{n})$ of $\operatorname{Vec}(X)$ which consists of all vector fields on ${\mathbb{A}}^{n}$ with constant divergence, see [Sha81, p. 191] and [FuKr18, Proposition 15.7.2]. For an algebraic subgroup $G\subseteq\operatorname{Aut}({\mathbb{A}}^{n})$, $\operatorname{Lie}(G)$ is a finite-dimensional Lie subalgebra of $\operatorname{Vec}^{\mathrm{c}}({\mathbb{A}}^{n})$.
Let $\mathfrak{G}$ be a connected ind-group and $\mathfrak{H}\subseteq\mathfrak{G}$ be a closed ind-subgroup. Then $\operatorname{Lie}(\mathfrak{H})$ is a Lie subalgebra of $\operatorname{Lie}(\mathfrak{G})$. However, it can happen (contrary to [Sha81, Theorem 1]) that $\operatorname{Lie}(\mathfrak{H})=\operatorname{Lie}(\mathfrak{G})$ while $\mathfrak{H}$ is a proper subgroup of $\mathfrak{G}$, see Example 9.5. In this respect, the ind-groups are more complex than the algebraic groups.
@@ 2.1. Algebraically generated subgroups of $\operatorname{Aut}(X)$ and Lie algebras
Let $G\subseteq\operatorname{Aut}(X)$ be a subgroup. One says that $G$ is algebraically generated if it is generated by a family $\{G_{i}\,|\,i\in I\}$ of connected algebraic subgroups of $\operatorname{Aut}(X)$. Throughout this subsection we assume that $G$ is algebraically generated. We associate to $G$ the Lie subalgebra
@@
generated by the Lie subalgebras $\operatorname{Lie}(G_{i})$. In fact, $L(G)$ depends only on $G$ and not on the choice of generating subgroups $G_{i}$, see [KrZa22, Theorem 2.3.1(4)].
For any ind-group $\mathfrak{G}$ we also define the canonical Lie algebra $L_{\mathfrak{G}}$, see [KrZa22, Sect. 7.1], where
@@
Consider also the Zariski closure $\overline{G}$ in the ind-group $\operatorname{Aut}(X)$. This is a closed ind-subgroup with Lie algebra $\operatorname{Lie}(\overline{G})$. Notice that $L(G)\subseteq L_{G}$ and the both are ideals in $\operatorname{Lie}\overline{G}$, see [KrZa22, Theorem 2.3.1 and Proposition 7.1.3].
Question 2.1 ([KrZa22, Question 1]).
@@
Then answer is affirmative provided $L$ is finite-dimensional.
Theorem 2.2 ([CD03, Theorem 1], [KrZa22, Theorem A]).
@@ 2.2. Locally finite endomorphisms and Lie subalgebras
Every locally finite endomorphism $\lambda$ has a uniquely defined additive Jordan decomposition $\lambda=\lambda_{s}+\lambda_{n}$ where $\lambda_{s}$ is semisimple, $\lambda_{n}$ locally nilpotent, and $\lambda_{s}\circ\lambda_{n}=\lambda_{n}\circ\lambda_{s}$.
@@
A subspace $L\subseteq\operatorname{Vec}(X)=\operatorname{Der}(\mathcal{O}(X))$ is called locally finite if for any $f\in\mathcal{O}(X)$ the subspace $Lf\subseteq\mathcal{O}(X)$ is finite dimensional.
If $L$ is locally finite, then it is finite dimensional, see [KrZa22, Lemma 1.6.2].
Let $L\subseteq\operatorname{Vec}(X)$ be a Lie subalgebra generated by a family of locally finite Lie subalgebras $L_{i}\subseteq\operatorname{Vec}(X)$, $i\in I$. There is the following question, see [KrZa22, Question 2].
Question 2.5.
@@
Corollary 2.6 ([KrZa22, Corollary of Theorem E]).
@@
Recall that a vector field on $X$ is locally nilpotent if so is the associated derivation $\partial\in\operatorname{Der}\mathcal{O}(X)$, that is, if for any $f\in\mathcal{O}(X)$ we have $\partial^{n}f=0$ for some $n=n(f)>0$. A weak form of Question 2.5 is as follows.
Question 2.7.
[…]
By Corollary 2.6 the latter is true if $\xi$ and $\eta$ define locally nilpotent derivations of $\mathcal{O}(X)$. Indeed, under the latter assumption we have by Theorem 2.2 $L=\operatorname{Lie}(G)$, where $G=\langle\exp(t\xi),\exp(t\eta)\rangle$ is an algebraic group.
3. Smoothness in ind-groups
By definition, an ind-variety $\mathfrak{V}$ comes equipped with a countable ascending filtration by algebraic varieties $\mathfrak{V}=\cup_{d}V_{d}$ where $V_{d}$ is closed in $V_{d+1}$. Another such filtration on $\mathfrak{V}$ is called admissible if the identity of $\mathfrak{V}$ defines isomorphism of the corresponding ind-varieties. There are different notions of smoothness of a point in an ind-variety.
@@ 3. Smoothness in ind-groups
  • $\mathfrak{V}$ is geometrically smooth in $x$ if there is an admissible filtration $\mathfrak{V}=\cup_{k}V_{d}$ such that $x\in V_{d}$ is a smooth point for all $d$.
@@
    Suppose that $\mathfrak{V}$ is connected and that $\mathfrak{W}$ has an admissible filtration $\mathfrak{W}=\cup_{k}W_{k}$ such that all $W_{k}$ are irreducible and normal. Then $\varphi$ is an isomorphism.
@@
    Suppose that $\mathfrak{V}$ is connected and $\mathfrak{W}$ is strongly smooth in a point $w\in\mathfrak{W}$. Then there is an open neighborhood $\mathfrak{V}^{\prime}$ of $\varphi^{-1}(w)$ in $\mathfrak{V}$ such that $\varphi|_{\mathfrak{V}^{\prime}}\colon\mathfrak{V}^{\prime}\to\mathfrak{W}$ is an open immersion. If $\mathfrak{W}$ is strongly smooth in every point, then $\varphi$ is an isomorphism.
@@
has a non-trivial kernel, see [BW00, Sect. 11, the last paragraph] and [FuKr18, Proposition 14.2.1]. In particular, $\varphi$ is not an isomorphism of ind-groups. In fact, there exists a unicuspidal curve $C$ in $\operatorname{Aut}({\mathbb{K}}\langle x,y\rangle)$ such that $\varphi(C)\subseteq\operatorname{Aut}({\mathbb{K}}[x,y])$ has a deeper cusp and so, is not isomorphic to $C$, see [FuKr18, Sect. 14.3].
@@
There exists as well a notion of algebraic smoothness of a point in an ind-variety. It has the following advantage: the point $e\in\mathfrak{G}$ is algebraically smooth in every affine ind-group $\mathfrak{G}$, see [Sha81, Theorem 3] and [Kum02, Theorem 4.3.7]. A geometrically smooth point of an ind-variety is algebraically smooth, see [Kum02, Lemma 4.3.4].
@@ 4. Nested ind-groups
The last assertion of Proposition 3.2 can be applied to nested ind-groups. Indeed, such a group is strongly smooth, see [FuKr18, Example 2.4.2]. Notice that a connected nested ind-group is algebraically generated, and any closed ind-subgroup of a nested ind-group also is nested.
Question 4.2.
Let $G\subseteq\operatorname{Aut}(X)$ be a connected nested ind-subgroup. Is it true that $G$ is closed?
See [Per23] for some partial results.
@@
If every finite subset of an ind-group $\mathfrak{G}$ (resp., of $\operatorname{Aut}(X)$) is contained in an algebraic group, then $\mathfrak{G}$ (resp., $\operatorname{Aut}(X)$) is nested.
Conjecture 4.4 ( [^1] Cf. [KPZ17, Proposition 2.10]. However, the proof of this proposition contains a mistake, as noted J.-P. Furter.).
Let $\mathfrak{G}=\bigcup_{i}G_{i}$ be a connected nested ind-subgroup of $\operatorname{Aut}(X)$, let $T\subseteq\mathfrak{G}$ be a maximal torus of $\mathfrak{G}$, and let $C^{0}(T)$ be the associate Cartan subalgebra, that is, the connected component of the centralizer of $T$. Then for any $g\in C^{0}(T)$ the conjugacy class of $g$ is closed in $\mathfrak{G}$.
@@
A subset $S$ of an ind-group $\mathfrak{G}$ is called weakly closed if for every algebraic subset $U\subseteq\mathfrak{G}$ such that $U\subseteq S$ we have $\overline{U}\subseteq S$.
@@
Let $g\in\operatorname{Aut}({\mathbb{A}}^{n})$ be locally finite. Does the weak closure of $C(g)$ contain the semisimple part of $g$?

### v2

@@ Abstract.
This is a brief overview of a few selected chapters on automorphism groups of affine varieties. It includes some open problems.
@@ 1. Introduction
We provide a brief overview of a few selected chapters on automorphism groups of affine varieties. It completes the existing literature on the subject, see e.g. [Arz23], [Fre17], [FuKr18], [Kal09], [KPZ17], [Kra89], [Kra96], [Kra17], [Lam24], [Miy24], [Sno89], [vdE00], [vdEKC21], [Wri17], etc.
Throughout the text, $X$ stands for an affine variety defined over an algebraically closed field ${\mathbb{K}}$ of characteristic zero, $\operatorname{Aut}(X)$ stands for the group of automorphisms of $X$, ${\mathbb{A}}^{n}$ stands for the affine $n$-space over ${\mathbb{K}}$ and $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$ resp. $\operatorname{{\mathbb{G}}_{\mathrm{m}}}$ stands for the additive resp. multiplicative group of the field ${\mathbb{K}}$.
If $\dim(X)=1$ then $\operatorname{Aut}(X)$ is an algebraic group. Moreover, any finite group can be realized as the full automorphism group of a smooth affine curve, see [Jel15, Proposition 7.2]. The properties of $\operatorname{Aut}(X)$ are ultimately related with the geometry of $X$. For instance, if $\operatorname{Aut}(X)$ is infinite, then $X$ is uniruled, i.e. covered by rational curves, see [Jel15, Theorem 1.1]. Already in the case of a smooth affine surface $X$, the group $\operatorname{Aut}(X)$ is often infinite-dimensional and has a rich algebraic structure, see the survey article [KPZ17]. The present survey is focused on various aspects of $\operatorname{Aut}(X)$ as an ind-group and on its Lie algebra. It contains a list of open problems partially borrowed in [FuKr18] and [KrZa24].
@@
[…]
The group of regular automorphisms $\operatorname{Aut}(X)$ is an affine ind-group, see e.g. [FuKr18, Theorem 5.1.1]. A closed algebraic subset of $A_{i}$ defines an algebraic family of automorphisms of $X$; such families were studied in [Ram64], see also [FuKr18, Section 3.3] and [Pop14].
A subgroup $G\subseteq\operatorname{Aut}(X)$ that is a closed algebraic subset of some $A_{i}$ has a natural structure of an algebraic group. It will be called an algebraic subgroup of $\operatorname{Aut}(X)$. Since $\operatorname{Aut}(X)$ is an affine ind-group, any algebraic subgroup of $\operatorname{Aut}(X)$ is an affine algebraic group. A faithful action of a linear algebraic group $G$ on $X$ defines an isomorphism of $G$ with an algebraic subgroup of $\operatorname{Aut}(X)$.
[…]
Let $\mathfrak{G}=\operatorname{Aut}(X)$ and let $\operatorname{Vec}(X)$ stand for the Lie algebra of regular vector fields on $X$. Recall that $\operatorname{Vec}(X)$ is naturally isomorphic to the Lie algebra $\operatorname{Der}(\mathcal{O}(X))$ of derivations of the ${\mathbb{K}}$-algebra $\mathcal{O}(X)$ of regular functions on $X$. In turn, $\operatorname{Lie}(\operatorname{Aut}(X))$ can be naturally identified with a subalgebra of $\operatorname{Vec}(X)$, see [FuKr18, Propositions 3.2.4 and 7.2.4]. For $X={\mathbb{A}}^{n}$ this is the subalgebra $\operatorname{Vec}^{\mathrm{c}}({\mathbb{A}}^{n})\subseteq\operatorname{Vec}({\mathbb{A}}^{n})$ of all vector fields on ${\mathbb{A}}^{n}$ with constant divergence, see [Sha81, p. 191] and [FuKr18, Proposition 15.7.2]. For an algebraic subgroup $G\subseteq\operatorname{Aut}(X)$, $\operatorname{Lie}(G)$ is a finite-dimensional Lie subalgebra of $\operatorname{Vec}(X)$.
Let $\mathfrak{G}$ be an ind-group and $\mathfrak{H}\subseteq\mathfrak{G}$ be a closed ind-subgroup. Then $\operatorname{Lie}(\mathfrak{H})$ is a Lie subalgebra of $\operatorname{Lie}(\mathfrak{G})$. However, it can happen (contrary to [Sha81, Theorem 1]) that $\operatorname{Lie}(\mathfrak{H})=\operatorname{Lie}(\mathfrak{G})$ while $\mathfrak{H}$ is a proper subgroup of a connected ind-group $\mathfrak{G}$, see Example 9.5. In this respect, the ind-groups are more complex than the algebraic groups.
@@ 2.1. Algebraically generated subgroups of $\operatorname{Aut}(X)$ and Lie algebras
Let $G\subseteq\operatorname{Aut}(X)$ be a subgroup. One says that $G$ is algebraically generated if it is generated by a family $\{G_{i}\,|\,i\in I\}$ of connected algebraic subgroups of $\operatorname{Aut}(X)$. Thus, an algebraically generated subgroup is connected.
Throughout this subsection we assume that $G$ is algebraically generated. We associate to $G$ the Lie subalgebra
@@
generated by the Lie subalgebras $\operatorname{Lie}(G_{i})$. In fact, $L(G)$ depends only on $G$ and not on the choice of generating subgroups $G_{i}$, see [KrZa24, Theorem 2.3.1(4)].
For any ind-group $\mathfrak{G}$ we also define the canonical Lie algebra $L_{\mathfrak{G}}$, see [KrZa24, Sect. 7.1], where
@@
Consider also the Zariski closure $\overline{G}$ in the ind-group $\operatorname{Aut}(X)$. This is a closed ind-subgroup with Lie algebra $\operatorname{Lie}(\overline{G})$. Notice that $L(G)\subseteq L_{G}$ and both are ideals in $\operatorname{Lie}\overline{G}$, see [KrZa24, Theorem 2.3.1 and Proposition 7.1.3].
Question 2.1 ([KrZa24, Question 1]).
@@
The answer is affirmative provided $L$ is finite-dimensional.
Theorem 2.2 ([CD03, Theorem 1], [KrZa24, Theorem A]).
@@ 2.2. Locally finite endomorphisms and Lie subalgebras
Every locally finite endomorphism $\lambda$ has a uniquely defined additive Jordan decomposition $\lambda=\lambda_{s}+\lambda_{n}$ where $\lambda_{s}$ is semisimple, $\lambda_{n}$ locally nilpotent and $\lambda_{s}\circ\lambda_{n}=\lambda_{n}\circ\lambda_{s}$. Recall that a locally finite endomorphism $\lambda_{s}$ is semisimple if the restriction of $\lambda_{s}$ to any finite-dimensional $\lambda_{s}$-invariant subspace of $V$ can be diagonalized.
@@
A subspace $L\subseteq\operatorname{Vec}(X)=\operatorname{Der}(\mathcal{O}(X))$ is called locally finite if any $f\in\mathcal{O}(X)$ is contained in a finite dimensional $L$-invariant subspace of $\mathcal{O}(X)$.
Every element of a locally finite subspace is locally finite as an endomorphism of $\mathcal{O}(X)$ and thus admits a Jordan decomposition.
Lemma 2.5 ([KrZa24, Lemma 1.6.2]).
A locally finite subspace $L\subseteq\operatorname{Vec}(X)$ is finite dimensional.
Let $L\subseteq\operatorname{Vec}(X)$ be a Lie subalgebra generated by a family of locally finite Lie subalgebras $L_{i}\subseteq\operatorname{Vec}(X)$, $i\in I$. There is the following question, see [KrZa24, Question 2].
Question 2.6.
@@
Corollary 2.7 ([KrZa24, Corollary of Theorem E]).
Consider a family $\{\eta_{i}\mid i\in I\}$ of locally nilpotent vector fields on $X$ and let $L:=\langle\eta_{i}\mid i\in I\rangle_{\text{\tiny Lie}}\subseteq\operatorname{Vec}(X)$ be the Lie algebra generated by the $\eta_{i}$. If $L$ is finite dimensional, then $L$ is locally finite.
Recall that a locally nilpotent vector field $\eta$ on $X$ defines a one-dimensional algebraic subgroup $U=\exp(t\eta)\subseteq\operatorname{Aut}(X)$ isomorphic to $\operatorname{{\mathbb{G}}_{\mathrm{a}}}$ such that $\operatorname{Lie}U={\mathbb{K}}\,\eta$.
A weak form of Question 2.6 is as follows.
Question 2.8.
Let $\xi,\eta\in\operatorname{Vec}(X)$ be locally finite. Assume that the Lie subalgebra $L=\langle\xi,\eta\rangle_{\text{\tiny Lie}}$ generated by $\xi$ and $\eta$ is finite dimensional. Does it follow that $L$ is locally finite?
By Corollary 2.7 the latter is true if $\xi$ and $\eta$ are locally nilpotent. Indeed, under the latter assumption we have by Theorem 2.2 $L=\operatorname{Lie}(G)$ where $G=\langle\exp(t\xi),\exp(t\eta)\rangle$ is an algebraic group.
3. Bijective morphisms of ind-groups
@@
  • $\mathfrak{V}$ is geometrically smooth in $x$ if there is an admissible filtration $\mathfrak{V}=\bigcup_{k}V_{d}$ such that $x\in V_{d}$ is a smooth point for all $d$.
@@
    Suppose that $\mathfrak{V}$ is connected and that $\mathfrak{W}$ has an admissible filtration $\mathfrak{W}=\bigcup_{k}W_{k}$ such that all $W_{k}$ are irreducible and normal. Then $\varphi$ is an isomorphism.
@@
    Suppose that $\mathfrak{V}$ is curve-connected and $\mathfrak{W}$ is strongly smooth in a point $w\in\mathfrak{W}$. Then there is an open neighborhood $\mathfrak{V}^{\prime}$ of $\varphi^{-1}(w)$ in $\mathfrak{V}$ such that $\varphi|_{\mathfrak{V}^{\prime}}\colon\mathfrak{V}^{\prime}\to\mathfrak{W}$ is an open immersion. If $\mathfrak{W}$ is strongly smooth in every point, then $\varphi$ is an isomorphism.
@@
has a non-trivial kernel, see [BW00, last paragraph of sect. 11] and [FuKr18, Proposition 14.2.1]. In particular, $\varphi$ is not an isomorphism of ind-groups. In fact, there exists a unicuspidal curve $C$ in $\operatorname{Aut}({\mathbb{K}}\langle x,y\rangle)$ such that $\varphi(C)\subseteq\operatorname{Aut}({\mathbb{K}}[x,y])$ has a deeper cusp and so, is not isomorphic to $C$, see [FuKr18, Sect. 14.3].
@@
It is not clear how the notions “strongly smooth” and “geometrically smooth” are related. There exists as well a notion of algebraic smoothness of a point in an ind-variety. It has the following advantage: the point $e\in\mathfrak{G}$ is algebraically smooth in every affine ind-group $\mathfrak{G}$, see [Sha81, Theorem 3] and [Kum02, Theorem 4.3.7]. A geometrically smooth point of an ind-variety is algebraically smooth, see [Kum02, Lemma 4.3.4].
@@ 4. Nested ind-groups
The last assertion of Proposition 3.2 can be applied to nested ind-groups. Indeed, such a group is strongly smooth, see [FuKr18, Example 2.4.2]. Notice that a connected nested ind-group is algebraically generated, and any closed ind-subgroup of a nested ind-group is nested.
Question 4.2 ([KrZa24, Question 4]).
Is it true that every connected nested ind-subgroup $G\subseteq\operatorname{Aut}(X)$ is closed in $\operatorname{Aut}(X)$?
The answer is affirmative, see [Per24, Theorem 6.6]. Furthermore, a subgroup of $\operatorname{Aut}(X)$ consisting of unipotent elements is nested if and only if it is closed in $\operatorname{Aut}(X)$, see [Per24, Theorem 8.11]. The proofs of these results exploit a new notion of a de Jonquières-like (dJ-like, for short) subgroup of $\operatorname{Aut}(X)$. The following analogue of the Lie–Kolchin theorem holds: every nested unipotent subgroup $U\subseteq\operatorname{Aut}(X)$ is contained in a dJ-like subgroup, see [Per24, Theorem 6.1]. Yet another ingredient in the proof of [Per24, Theorem 6.6] is the following version of the Levi decomposition for nested subgroups, see [KPZ17, Theorem 2.11 and Corollary 2.12].
Theorem 4.3.
[…]
There is the following conjectural criterion of nestedness.
@@
Assume that every finite subset of an ind-group $\mathfrak{G}$ is contained in an algebraic group. Then $\mathfrak{G}$ is nested.
@@
A subset $S$ of an ind-group $\mathfrak{G}$ is called weakly closed if for every algebraic subset $U\subseteq\mathfrak{G}$ such that $U\subseteq S$ we have $\overline{U}\subseteq S$. The union $\cup_{U}\overline{U}$ for all $U$ as before is called the weak closure of $S$.
@@
Let $g\in\operatorname{Aut}({\mathbb{A}}^{n})$ be locally finite. Does the (weak) closure of $C(g)$ contain the semisimple part of $g$?